# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

### v2

@@ 6 Conclusion
In this paper, we identify and formalize modality interference as a concrete manifestation of the broader cross-modality competency problem in Multimodal Large Language Models—namely, the inability to distinguish task-relevant from irrelevant modality signals. Through a designed perturbation-based causal evaluation experiment, we demonstrate that even state-of-the-art MLLMs systematically exhibit degraded performance under irrelevant but misleading inputs, revealing a fundamental vulnerability in their inference-time reasoning.
To mitigate this issue, we propose a robust fine-tuning strategy that combines modality-specific data augmentation, consistency regularization, and adversarial perturbation in the embedding space. These interventions explicitly constrain the model to produce stable outputs under spurious modality shifts, thereby reducing reliance on non-causal correlations and improving robustness. Extensive experiments across diverse architectures, scales, and task regimes confirm that our approach consistently improves both unimodal reasoning and multimodal generalization. Our findings suggest that enhancing cross-modality competency requires not only better alignment, but also causal safeguards against shortcut learning from irrelevant modalities.
@@ Appendix A Appendix Summary
Finetuning StrategiesAppendix˜B): We elaborate on our design choice to freeze the Q-Former in InstructBLIP-based models. This decision is motivated by the need to retain strong visual representations while avoiding overfitting to perturbed or misleading multimodal inputs. ( Table˜12 records the size for each dataset)
Detailed Experimental ResultsAppendix˜C): This section includes three key tables— Table˜4,  Table˜5 and  Table˜6—which report model performance on unimodal and multimodal tasks under various perturbation settings and ablation conditions(additional models included).  Table˜7 records the performance of different vanilla MLLMs under modality interference across modality-heavy datasets. We also include radar plots ( Figure˜4) that visualize task-wise robustness across different MLLMs. We provided the detailed experimental results on Qwen2.5-VL-7b [3] and InstructBlip-Vicuna-13b [2] and make further discussion on the selection of specific consistency loss. In  Table˜8, we examine the generalization benefits of adversarial training by evaluating robustness under two types of out-of-distribution (OOD) perturbations: real-world OCR noise (from FUNSD [39]) and unrelated screenshots (from RICO [40]). In  Table˜9, we assess the impact of Chain-of-Thought prompting in mitigating modality interference, comparing its effectiveness against our method and standard baselines across both visual and textual modalities. In  Table˜10, we report results on the free-form generative VQA benchmark TextVQA [41], highlighting our method’s generalizability beyond multiple-choice formats.
Hyperparameter SettingsAppendix˜D): We present full training configurations used in our experiments, including optimization strategies, perturbation settings, and sampling ratios for different task types. This section enables reproducibility and highlights the computational efficiency of our proposed training scheme. We provide parameter analysis on iterations of PGD in  Figure˜5.
Compute Resource DetailsAppendix˜E): We document hardware specifications, training durations, and resource costs for models of different scales. These details contextualize the feasibility of our approach in academic environments.
LimitationsAppendix˜F): We discuss the granularity of our current modality interference analysis, the selections of perturbations, and propose directions for more fine-grained future studies.
Broader ImpactsAppendix˜G): Finally, we reflect on the ethical implications and societal benefits of our research. While our methods improve model robustness and alignment, we also acknowledge the dual-use nature of adversarial perturbations and advocate for safety-aware deployment.
@@ Appendix C Detailed Experimental Results
Please see  Table˜4,  Table˜5,  Figure˜4,  Table˜7,  Table˜6,  Table˜8,  Table˜9 and  Table˜10 for more details.
@@
Qwen2.5-VL Models [3]
Qwen2.5-VL-3B [3] Vanilla 63.0 (100%) 81.8 (100%) 72.3 (100%) 72.5 (100%)
@@
Instructblip-Vicuna Models [2]
InstructBlip-7B [2] Vanilla 52.1 (100%) 65.5 (100%) 54.8 (100%) 56.4 (100%)
@@
LLaVA1.5 Models [14]
LLaVA-1.5-7B [14] Vanilla 64.5 (100%) 64.3 (100%) 63.4 (100%) 63.8 (100%)
@@
Note. While Qwen2.5-VL was originally instruction-tuned with proprietary in-house data [3], our reproduced version uses only publicly available LLaVA instruction-tuning data. Even under this constraint and without access to VQA-specific tuning samples, our models achieve comparable or even better performance across all VQA datasets—highlighting the robustness and effectiveness of our proposed perturbation-consistent fine-tuning strategy.
Table 7: Performance (%) of Vinilla models under modality interference across four datasets. We show accuracy under clean (origin) and various perturbations: Left: Mini-ImageNet and Caltech-101; Right: OpenBookQA and MMLU. UF = Unrelated Facts, MD = Misleading Descriptions, RP = Random Pixels, RI = Real Image, FB & FW = Full Black/White Canvas.
@@
InternVL2-2B [42] 91.9 88.6 25.5 94.6 91.3 33.2 46.3 36.2 45.2 45.3 39.1 36.9 39.3 39.2
LLaVA-1.5-7B [14] 95.3 93.4 43.5 97.0 95.9 57.4 62.4 56.4 62.5 63.4 46.3 45.2 45.9 45.8
LLaVA-Next-7B [14] 93.4 90.0 28.5 97.0 93.4 31.6 54.6 52.9 55.9 55.5 45.9 45.0 45.9 45.8
LLaVA-1.5-13B [14] 95.6 94.1 73.0 97.9 97.1 77.4 65.9 63.8 68.0 69.1 51.8 50.8 52.7 52.7
InstructBlip-7B [2] 92.0 87.1 13.6 90.3 90.2 17.5 50.8 46.2 50.9 50.7 36.3 35.2 36.2 36.7
InstructBlip-13B [2] 93.0 81.6 50.9 94.1 82.7 51.0 44.7 39.4 45.6 46.1 40.1 37.4 40.5 42.9
QwenVL2-2B [43] 98.7 98.6 75.8 99.1 99.1 66.8 59.6 54.7 63.4 63.8 49.5 44.6 49.7 50.0
QwenVL2.5-3B [3] 98.9 98.5 94.9 98.8 99.0 94.4 79.9 74.6 80.0 79.7 63.5 61.1 64.0 63.6
QwenVL2-7B [43] 99.0 99.1 96.3 99.6 99.5 97.6 82.5 80.9 83.7 83.2 66.9 65.3 67.7 67.8
QwenVL2.5-7B [3] 99.3 99.3 96.5 99.3 99.4 96.8 86.2 80.9 86.3 86.4 69.5 67.6 69.1 69.1
@@ C.1 Perturbation-based Evaluation Experiment Results
We conduct a controlled perturbation-based evaluation across various MLLMs, as shown in Table˜7. Our results reveal that both vision and language tasks are vulnerable to cross-modal interference. In vision classification tasks, misleading textual descriptions (e.g., text contradicting image content) lead to severe performance drops. For example, InternVL2-2B and InstructBLIP-7B on Mini-ImageNet drop from 91.9% to 25.5% and from 92.0% to 13.6%, respectively. Conversely, for language tasks such as OpenBookQA and MMLU, irrelevant visual inputs—particularly semantically unrelated real images—also degrade performance. LLaVA-1.5-7B drops from 46.3% to 45.2% on MMLU, while InstructBLIP-13B sees over 5 points of degradation.
@@
To enable a more equitable comparison with existing multimodal models, we extend our method to two additional backbones: InstructBLIP-Vicuna-13B and QwenVL-2.5-7B. As shown in Table˜6, our approach consistently improves performance across multiple VQA benchmarks, even under different instruction-tuning conditions, demonstrating its robustness and general applicability.
@@ C.3 Discussion on KL&JS use for Consistency Regularization
Although both KL and JS divergence serve as effective objectives for consistency regularization, we find that JS consistently achieves slightly better results across most settings. Specifically, in both unimodal tasks (e.g., Mini-ImageNet, MMLU) and multimodal reasoning benchmarks (e.g., ScienceQA, SeedBench), JS-regularized models consistently outperform their KL counterparts by a small but observable margin. This trend holds across different model backbones and training configurations, including our final unified method (see “Ours” rows in Table 4 and 6). This suggests a marginal advantage of JS regularization in enhancing model robustness.
C.4 Improving Generalizability through Adversarial Training
  • Document OCR noise: Real-world noisy OCR snippets are sampled from the FUNSD dataset [39] and inserted as irrelevant textual distractors into visual classification tasks (Mini-ImageNet, Caltech-101).
  • Unrelated screenshots: Unrelated UI screenshots are drawn from the RICO dataset [40] and added as visual distractors to language-dominant VQA tasks (OpenBookQA, MMLU).
Each experiment is repeated multiple times, and we report average accuracy across runs (standard deviation << 0.1). Results are presented in Table 8. Across all model scales and task types, adversarial training consistently improves robustness to both types of perturbations. These findings indicate that the additional training overhead introduced by PGD is well-justified by the improved generalization to unseen distribution shifts—a desirable property for reliable deployment in real-world settings.
Setting Model Mini-ImageNet Caltech-101 OpenBookQA MMLU
Origin OCR Origin OCR RandPixels Screenshot RandPixels Screenshot
SFT instructblip-vicuna-7b 98.5 95.7 99.0 98.4 75.0 74.4 50.0 49.2
SFT + PGD instructblip-vicuna-7b 98.7 97.5 99.5 98.8 76.8 76.2 49.3 49.3
Ours instructblip-vicuna-7b 98.4 97.3 99.2 99.0 79.0 77.8 50.2 49.2
vanilla llava-1.5-7b 95.3 88.9 97.0 92.8 62.4 50.2 46.3 44.8
SFT llava-1.5-7b 98.2 98.0 98.5 98.0 78.6 78.0 51.1 50.6
SFT + PGD llava-1.5-7b 98.7 98.2 98.7 98.4 81.7 81.3 50.6 51.3
Ours llava-1.5-7b 98.6 98.2 99.3 99.0 81.8 81.2 51.5 51.3
vanilla qwen2.5-vl-7b 99.3 99.2 99.1 99.2 85.9 69.3 69.3 57.0
SFT qwen2.5-vl-7b 99.6 99.5 99.7 99.7 90.2 85.8 70.4 63.7
SFT + PGD qwen2.5-vl-7b 99.4 99.5 99.6 99.6 91.7 91.5 70.4 69.7
Ours qwen2.5-vl-7b 99.6 99.5 99.6 99.5 90.9 90.7 69.8 69.7
C.5 Evaluating Chain-of-Thought Prompting for Modality Interference Mitigation
To further investigate the potential of prompt-based methods in mitigating modality interference, we conduct additional experiments using Chain-of-Thought [44] style prompting. This approach aims to encourage structured reasoning by guiding the model through an explicit reasoning process before producing its final answer.
@@ 3. Are there any misleading parts?
Model performance is evaluated under both clean inputs and perturbed conditions. For each task, we report accuracy on the original input as well as on the strongest-performing perturbation type (e.g., OCR insertion for visual classification tasks, or unrelated screenshots for VQA and language tasks). Results are presented in Table 9.
@@
C.6 Evaluation on Free-Form VQA
To further assess the generalizability of our method beyond multiple-choice VQA tasks, we evaluate on TextVQA [41], a free-form generative visual question answering dataset that requires reasoning over both textual and visual content in natural images. We follow the benchmark adopted by LLaVA [1], which evaluates a model’s ability to both recognize textual characters within images and effectively handle noisy outputs generated by OCR systems.

### v3

@@ 6 Conclusion
In this paper, we identify and formalize modality interference as a concrete manifestation of the broader cross-modality competency problem in Multimodal Large Language Models—namely, the inability to distinguish task-relevant from irrelevant modality signals. Through a designed perturbation-based causal evaluation experiment, we demonstrate that even state-of-the-art MLLMs systematically exhibit degraded performance under irrelevant but misleading inputs, revealing a fundamental vulnerability in their inference-time reasoning. To mitigate this issue, we propose a robust fine-tuning strategy that combines modality-specific data augmentation, consistency regularization, and adversarial perturbation in the embedding space. These designs explicitly constrain the model to produce stable outputs under spurious modality shifts, thereby reducing reliance on non-causal correlations and improving robustness. Extensive experiments across diverse architectures, scales, and task regimes confirm that our approach consistently improves both unimodal reasoning and multimodal generalization, achieving Pareto-optimal performance.
Ethics Statement
Our research focuses on analyzing and mitigating modality interference in Multimodal Large Language Models. All experiments are conducted on publicly available benchmark datasets, including Mini-ImageNet, Caltech-101, OpenBookQA, MMLU, ScienceQA, MM-Bench, and Seed-Bench, which contain no personally identifiable or sensitive information beyond what is publicly released. We do not foresee direct risks of harm to individuals or groups arising from this research. Nevertheless, potential societal impacts include bias amplification or misinterpretation when deploying MLLMs in real-world applications. We note these risks and emphasize that our contributions are methodological and diagnostic rather than application-specific. No human subjects were involved, and no IRB approval was required. All funding sources are acknowledged in the main paper. Further discussions of limitations, broader impacts and LLM use are provided in Appendix Appendix F, Appendix Appendix G and Appendix Appendix H.
Reproducibility Statement
We are committed to ensuring the reproducibility of our work.
  • Code and Implementation: We will release a full open-source codebase, including data processing, training, and evaluation scripts, upon publication.
  • Datasets: All datasets used in this work are publicly available (Mini-ImageNet, Caltech-101, OpenBookQA, MMLU, ScienceQA, MM-Bench, Seed-Bench). Detailed preprocessing steps and dataset conversions into unified multiple-choice VQA format are described in  Section 5 and Appendix Appendix C.
  • Model and Training Details: Hyperparameters (learning rates, batch sizes, epochs, optimizer choices) and architectural specifications are reported in  Section 5 and Appendix Appendix B, Appendix D, Appendix E.
  • Evaluation: Metrics, baselines, and evaluation protocols are fully documented in  Section 5 and Appendix Appendix C with complete ablation results.
Together, these materials should enable independent researchers to reproduce our findings.
@@ Appendix A Appendix Summary
Finetuning Strategies (Appendix B): We elaborate on our design choice to freeze the Q-Former in InstructBLIP-based models. This decision is motivated by the need to retain strong visual representations while avoiding overfitting to perturbed or misleading multimodal inputs. ( Table 13 records the size for each dataset)
Detailed Experimental Results (Appendix C): This section includes three key tables— Table 5,  Table 6 and  Table 7—which report model performance on unimodal and multimodal tasks under various perturbation settings and ablation conditions(additional models included).  Table 8 records the performance of different vanilla MLLMs under modality interference across modality-heavy datasets. We also include radar plots ( Figure 4) that visualize task-wise robustness across different MLLMs. We provided the detailed experimental results on Qwen2.5-VL-7b (Bai et al., 2025) and InstructBlip-Vicuna-13b (Luo et al., 2023) and make further discussion on the selection of specific consistency loss. In  Table 11, we examine the generalization benefits of adversarial training by evaluating robustness under two types of out-of-distribution (OOD) perturbations: real-world OCR noise (from FUNSD (Jaume et al., 2019)) and unrelated screenshots (from RICO (Deka et al., 2017)). In  Table 9, we assess the impact of Chain-of-Thought prompting in mitigating modality interference, comparing its effectiveness against our method and standard baselines across both visual and textual modalities. In  Table 10, we report results on the free-form generative VQA benchmark TextVQA (Singh et al., 2019), highlighting our method’s generalizability beyond multiple-choice formats.
Hyperparameter Settings (Appendix D): We present full training configurations used in our experiments, including optimization strategies, perturbation settings, and sampling ratios for different task types. This section enables reproducibility and highlights the computational efficiency of our proposed training scheme. We provide parameter analysis on iterations of adversarial training in  Figure 5.
Compute Resource Details (Appendix E): We document hardware specifications, training durations, and resource costs for models of different scales. These details contextualize the feasibility of our approach in academic environments.
Limitations (Appendix F): We discuss the granularity of our current modality interference analysis, the selections of perturbations, and propose directions for more fine-grained future studies.
Broader Impacts (Appendix G): We reflect on the ethical implications and societal benefits of our research. While our methods improve model robustness and alignment, we also acknowledge the dual-use nature of adversarial perturbations and advocate for safety-aware deployment.
LLM Use (Appendix H): Finally, we clarify that LLMs were only used to polish the writing of this paper.
@@ Appendix C Detailed Experimental Results
Please see  Table 5,  Table 6,  Figure 4,  Table 8,  Table 7,  Table 11,  Table 9 and  Table 10 for more details.
@@
Qwen2.5-VL Models (Bai et al., 2025)
Qwen2.5-VL-3B Bai et al. (2025) Vanilla 63.0 (100%) 81.8 (100%) 72.3 (100%) 72.5 (100%)
@@
Instructblip-Vicuna Models (Luo et al., 2023)
InstructBlip-7B Luo et al. (2023) Vanilla 52.1 (100%) 65.5 (100%) 54.8 (100%) 56.4 (100%)
@@
LLaVA1.5 Models (Liu et al., 2023a)
LLaVA-1.5-7B Liu et al. (2023a) Vanilla 64.5 (100%) 64.3 (100%) 63.4 (100%) 63.8 (100%)
@@
Note. While Qwen2.5-VL was originally instruction-tuned with proprietary in-house data (Bai et al., 2025), our reproduced version uses only publicly available LLaVA instruction-tuning data. Even under this constraint and without access to VQA-specific tuning samples, our models achieve comparable or even better performance across all VQA datasets—highlighting the robustness and effectiveness of our proposed perturbation-consistent fine-tuning strategy.
Table 8: Performance (%) of Vinilla models under modality interference across four datasets. We show accuracy under clean (origin) and various perturbations: Left: Mini-ImageNet and Caltech-101; Right: OpenBookQA and MMLU. UF = Unrelated Facts, MD = Misleading Descriptions, RP = Random Pixels, RI = Real Image, FB & FW = Full Black/White Canvas. (The results are averaged on multiple runs with standard deviation << 0.2)
@@
InternVL2-2B (Chen et al., 2024b) 91.9 88.6 25.5 94.6 91.3 33.2 46.3 36.2 45.2 45.3 39.1 36.9 39.3 39.2
LLaVA-1.5-7B (Liu et al., 2023a) 95.3 93.4 43.5 97.0 95.9 57.4 62.4 56.4 62.5 63.4 46.3 45.2 45.9 45.8
LLaVA-Next-7B (Liu et al., 2024) 93.4 90.0 28.5 97.0 93.4 31.6 54.6 52.9 55.9 55.5 45.9 45.0 45.9 45.8
LLaVA-Next-34B (Liu et al., 2024) 98.0 96.3 90.2 99.1 97.3 93.6 87.7 85.7 88.4 88.0 71.5 70.9 71.5 71.5
LLaVA-Next-72B (Liu et al., 2024) 97.7 97.8 83.0 98.9 98.8 91.2 88.2 86.2 89.1 88.9 73.6 72.9 73.9 73.9
LLaVA-Next-110B (Liu et al., 2024) 98.4 98.3 93.3 98.4 98.3 93.1 89.5 89.2 89.9 89.7 73.5 73.0 73.9 73.9
LLaVA-1.5-13B (Liu et al., 2023a) 95.6 94.1 73.0 97.9 97.1 77.4 65.9 63.8 68.0 69.1 51.8 50.8 52.7 52.7
InstructBlip-7B (Luo et al., 2023) 92.0 87.1 13.6 90.3 90.2 17.5 50.8 46.2 50.9 50.7 36.3 35.2 36.2 36.7
InstructBlip-13B (Luo et al., 2023) 93.0 81.6 50.9 94.1 82.7 51.0 44.7 39.4 45.6 46.1 40.1 37.4 40.5 42.9
QwenVL2-2B (Wang et al., 2024b) 98.7 98.6 75.8 99.1 99.1 66.8 59.6 54.7 63.4 63.8 49.5 44.6 49.7 50.0
QwenVL2.5-3B (Bai et al., 2025) 98.9 98.5 94.9 98.8 99.0 94.4 79.9 74.6 80.0 79.7 63.5 61.1 64.0 63.6
QwenVL2-7B (Wang et al., 2024b) 99.0 99.1 96.3 99.6 99.5 97.6 82.5 80.9 83.7 83.2 66.9 65.3 67.7 67.8
QwenVL2.5-7B (Bai et al., 2025) 99.3 99.3 96.5 99.3 99.4 96.8 86.2 80.9 86.3 86.4 69.5 67.6 69.1 69.1
@@ C.1 Perturbation-based Evaluation Experiment Results
We conduct a controlled perturbation-based evaluation across various MLLMs, as shown in Table 8. Our results reveal that both vision and language tasks are vulnerable to cross-modal interference. In vision classification tasks, misleading textual descriptions (e.g., text contradicting image content) lead to severe performance drops. For example, InternVL2-2B and InstructBLIP-7B on Mini-ImageNet drop from 91.9% to 25.5% and from 92.0% to 13.6%, respectively. Conversely, for language tasks such as OpenBookQA and MMLU, irrelevant visual inputs—particularly semantically unrelated real images—also degrade performance. LLaVA-1.5-7B drops from 46.3% to 45.2% on MMLU, while InstructBLIP-13B sees over 5 points of degradation.
@@
We further observe a clear scaling trend within the LLaVA-Next family. As model size increases from 7B to 34B, 72B, and 110B, performance under perturbations steadily improves, reflecting stronger representation power and enhanced robustness to spurious cues. For instance, LLaVA-Next-7B achieves 90.3% on Mini-ImageNet (Orig) but drops to 28.5% with misleading descriptions, whereas LLaVA-Next-110B maintains 98.4% and 93.3% under the same conditions. Similarly, on MMLU, accuracy under irrelevant real images increases from 45.9% (7B) to 73.6% (72B). These results confirm that scaling up helps mitigate modality interference. However, such gains come at substantial computational and resource costs, and the improvements remain incremental relative to the clean–perturbed gap. This underscores that scaling alone is insufficient, and more targeted interventions—such as our proposed framework—are necessary for robust cross-modal reasoning.
@@
To enable a more equitable comparison with existing multimodal models, we extend our method to two additional backbones: InstructBLIP-Vicuna-13B and QwenVL-2.5-7B. As shown in Table 7, our approach consistently improves performance across multiple VQA benchmarks, even under different instruction-tuning conditions, demonstrating its robustness and general applicability.
@@ C.3 Discussion on KL&JS use for Consistency Regularization
Although both KL and JS divergence serve as effective objectives for consistency regularization, we find that JS consistently achieves slightly better results across most settings. Specifically, in both unimodal tasks (e.g., Mini-ImageNet, MMLU) and multimodal reasoning benchmarks (e.g., ScienceQA, SeedBench), JS-regularized models consistently outperform their KL counterparts by a small but observable margin. This trend holds across different model backbones and training configurations, including our final unified method (see “Ours” rows in Table 5 and 7). This suggests a marginal advantage of JS regularization in enhancing model robustness.
C.4 Evaluating Chain-of-Thought Prompting for Modality Interference Mitigation
To further investigate the potential of prompt-based methods in mitigating modality interference, we conduct additional experiments using Chain-of-Thought Wei et al. (2022) style prompting. This approach aims to encourage structured reasoning by guiding the model through an explicit reasoning process before producing its final answer.
@@ 3. Are there any misleading parts?
Results are presented in Table 9. The results suggests that structured reasoning alone cannot resolve the interference problem, as the issue stems from misaligned cross-modal representations rather than shallow reasoning steps.
@@
C.5 Evaluation on Free-Form VQA
To further assess the generalizability of our method beyond multiple-choice VQA tasks, we evaluate on TextVQA (Singh et al., 2019), a free-form generative visual question answering dataset that requires reasoning over both textual and visual content in natural images. We follow the benchmark adopted by LLaVA Liu et al. (2023b), which evaluates a model’s ability to both recognize textual characters within images and effectively handle noisy outputs generated by OCR systems.
@@ C.5 Evaluation on Free-Form VQA
C.6 Improving Generalizability through Adversarial Training
  • Document OCR noise: Real-world noisy OCR snippets are sampled from the FUNSD dataset (Jaume et al., 2019) and inserted as irrelevant textual distractors into visual classification tasks (Mini-ImageNet, Caltech-101).
  • Unrelated screenshots: Unrelated UI screenshots are drawn from the RICO dataset (Deka et al., 2017) and added as visual distractors to language-dominant VQA tasks (OpenBookQA, MMLU).
Each experiment is repeated multiple times, and we report average accuracy across runs (standard deviation << 0.1). Results are presented in Table 11. Across all model scales and task types, adversarial training consistently improves robustness to both types of perturbations. These findings indicate that the additional training overhead introduced by adversarial perturbation is well-justified by the improved generalization to unseen distribution shifts—a desirable property for reliable deployment in real-world settings.
Setting Model Mini-ImageNet Caltech-101 OpenBookQA MMLU
Origin OCR Origin OCR RandPixels Screenshot RandPixels Screenshot
SFT instructblip-vicuna-7b 98.5 95.7 99.0 98.4 75.0 74.4 50.0 49.2
SFT + ADV instructblip-vicuna-7b 98.7 97.5 99.5 98.8 76.8 76.2 49.3 49.3
Ours instructblip-vicuna-7b 98.4 97.3 99.2 99.0 79.0 77.8 50.2 49.2
vanilla llava-1.5-7b 95.3 88.9 97.0 92.8 62.4 50.2 46.3 44.8
SFT llava-1.5-7b 98.2 98.0 98.5 98.0 78.6 78.0 51.1 50.6
SFT + ADV llava-1.5-7b 98.7 98.2 98.7 98.4 81.7 81.3 50.6 51.3
Ours llava-1.5-7b 98.6 98.2 99.3 99.0 81.8 81.2 51.5 51.3
vanilla qwen2.5-vl-7b 99.3 99.2 99.1 99.2 85.9 69.3 69.3 57.0
SFT qwen2.5-vl-7b 99.6 99.5 99.7 99.3 90.2 85.8 70.4 63.7
SFT + ADV qwen2.5-vl-7b 99.4 99.5 99.6 99.6 91.7 91.5 70.4 69.7
Ours qwen2.5-vl-7b 99.6 99.5 99.6 99.5 90.9 90.7 69.8 69.7